OpenStudy (anonymous):

Find the first five non-zero terms of power series representation centered at x=0 for the function below. f(x)=7/(1−x^2)

OpenStudy (anonymous):

I found the formula, I'm just really confused on finding the first five terms

OpenStudy (anonymous):

7 * series from n=0 to infinity of x^(2n)

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

the first thing i would do is factor out 7

OpenStudy (anonymous):

I think I did...? idk

OpenStudy (anonymous):

$$ \Large { \frac{7}{1-x^2}= 7\cdot \frac{1}{1-x^2}= 7 \sum_{k=0}^{n}x^{2k} } $$

OpenStudy (anonymous):

OH yup! That's what I got!

OpenStudy (anonymous):

Okay, I realized I just left the n.

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (anonymous):

$$ \Large { \frac{7}{1-x^2}\\= 7\cdot \frac{1}{1-x^2}= 7 \sum_{k=0}^{n}x^{2k} \\ = 7 ~(x^{2 \cdot 0 } + x^{2\cdot 1} + x^{2\cdot 2} + x^{2\cdot 3}+x^{2\cdot 4}... ) \\ = 7 ~(1 + x^2 + x^4 + x^{6}+x^8... ) \\ = 7 + 7x^2 + 7x^4 + 7x^6 +7x^8... \\ } $$

OpenStudy (anonymous):

your welcome